OpenStudy (anonymous):

Miranda wants to give her 14-year-old daughter $20,000 when she turns 18. How much does she need to put in the bank now if the interest rate is 10 percent per year? $12,418.43 $13,660.27 $15,026.30

OpenStudy (anonymous):

compound or simple interest?

OpenStudy (anonymous):

i dont know it just gives me this fomula p*(1+i)^t

OpenStudy (anonymous):

Hey, Wolf? Do you know if i=10 or .1?

OpenStudy (anonymous):

p*(1+10)^4=20,000, right?

OpenStudy (wolf1728):

The formula you will need is: Principal = Total ÷ (1 + Rate)^ Years Prin = 20,000 / (1.10)^4 Prin = 20,000 / 1.4641 Principal = 13,660.27 The formula is on this page: http://1728.org/compint2.htm and a calculator to check your work is on this page: http://1728.org/compint.htm

OpenStudy (wolf1728):

Hi raven it is .1

OpenStudy (anonymous):

Oh, okay. Thank you for the help. :) Have a great night.

OpenStudy (wolf1728):

Glad to have helped out. Yes, I'll get going for now.